# DreamRunner: Fine-Grained Compositional Story-to-Video Generation with Retrieval-Augmented Motion Adaptation

### v3

@@ 3.3 Spatial-Temporal-Region-Based Diffusion
We build our model on CogVideoX [68], a text-to-video generation model designed on top of a Diffusion Transformer (DiT). Unlike methods that use separate spatial and temporal attention for efficient video modeling, CogVideoX employs a 3D full attention module, integrating self-attention across concatenated embeddings of all visual latents and the text condition embeddings. We extend this module to enable region-specific conditioning via masking, aligning different regions with their respective text descriptions. Specifically, given a fine-grained plan with $N$ region-specific text descriptions $C_{1},C_{2},\dots,C_{N}$ and corresponding layouts $L_{1},L_{2},\dots,L_{N}$ across frames, we encode each text condition $C_{i}$ to produce embeddings $T_{1},T_{2},\dots,T_{N}$ (Fig. 2 top right). At each attention layer, we identify the visual latents corresponding to each layout $L_{i}$ in the latent space. We then perform masked self-attention on the concatenation of $T_{1},T_{2},\dots,T_{N}$ and $L_{1},L_{2},\dots,L_{N}$. The self-attention mask is defined as follows: for each region’s visual latents $L_{i}$, attention is allowed to its corresponding text condition embeddings $T_{i}$ and all visual latents $L_{1},L_{2},\dots,L_{N}$. Conversely, for each condition embeddings $T_{i}$, attention is restricted to itself and its corresponding latents $L_{i}$. This design ensures each region is conditioned on its specific textual description while maintaining interactions among visual latents through unmasked attention among $L_{1},L_{2},\dots,L_{N}$. No modifications are made to other modules in the base model, preserving the integrity of its original architecture. A visualization example of such masking strategy is contained in Fig. 3.
@@
4 Experiments
In this section, we first introduce the evaluation datasets and evaluation metrics details in Sec. 4.1, then compare our DreamRunner with prior methods on story-to-video generation in Sec. 4.2. Next, we present detailed ablation studies on the necessity of RAG and effectiveness of SR3AI in Sec. 4.3, and demonstrate the generalizability of our DreamRunner to improve compositional text-to-video generation on T2V-CompBench [54] in Sec. 4.5. Moreover, we show the effectiveness of RAG for learning the motion prior on a more comprehensive motion dataset we collect in Sec. 4.6. Lastly, we present qualitative comparison between our DreamRunner and previous approaches in Sec. 4.4.
4.1 Experimental Setups
Evaluation Datasets. We evaluate DreamRunner on two tasks: (1) story-to-video generation, and (2) compositional text-to-video generation. The first task focuses on the model’s ability to follow the text closely while maintaining character and scene consistency throughout the story. The second task assesses various aspects of compositionality in video generation. For (1) story-to-video generation, we collect and introduce a new benchmark dataset, DreamStorySet. Specifically, we collect 10 characters, including 6 from existing customization datasets (CustomConcept101 [30] and Dreambooth [50]), and 4 with generation models (FLUX [4]). (featuring two motions per scene) and three multi-character stories (featuring two or three motions per scene). Each story comprises 5 to 8 scenes, incorporating a total of 64 diverse motions throughout. We focus on single-character stories for quantitative evaluation of SVG models and reserve multi-character stories for qualitative evaluation. For (2) compositional text-to-video generation, we use the T2V-CompBench [54] to benchmark the performance of DreamRunner, where we select six dimensions except numeracy.
Evaluation Metrics. We evaluate our generated storytelling videos across multiple dimensions, including Character Consistency (Frame-Reference-Image CLIP/DINO scores), Full-Narration-Text Alignment per Scene (Image/Video-Text CLIP/ViCLIP scores), Fine-Grained-Text Alignment per Scene (Image/Video-Text CLIP/ViCLIP scores), Transition (Frame-Frame DINO score), Visual Quality (LAION aesthetic quality score [51]/MUSIQ Imaging quality score [27]/and AMT video smoothness score [34]). Detailed descriptions of the metrics and computation methods can be found in the appendix. We follow similar metrics to T2V-ComBench [54] for evaluating compositional text-to-video generation.
4.2 Story-To-Video Generation Evaluation
We compare DreamRunner with previous SoTAs (VideoDirectorGPT [37] and VLogger [81]) on our DreamStorySet dataset for story-to-video generation. For fair comparison, we decouple each scene narration into two single-motion descriptions, generating separate videos that are then combined into a single-scene video.
As shown in Table 1, DreamRunner improves CLIP/DINO scores by 13.1%/33.4% over VLogger, demonstrating the effectiveness of our learned subject prior and region-based LoRA injection for character consistency.
To evaluate text-following capability, we assess both full-prompt adherence and fine-grained event alignment. DreamRunner improves CLIP/iCLIP scores consistently on both settings, showing superior alignment with both full-scene and fine-grained event descriptions.
For transition quality, we compute the DINO-based transition score to measure scene and event consistency. DreamRunner significantly improves transitions by 27.2% over VLogger, highlighting the effectiveness of SR3AI in generating sequential events in a single scene.
Lastly, we evaluate visual quality across aesthetic quality, imaging quality, and motion smoothness. DreamRunner significantly enhances aesthetics while slightly improving the other two, demonstrating its capability to generate high-quality videos that adhere to complex scene descriptions while ensuring smooth event transitions.
Figure 4: Qualitative comparison and ablations of DreamRunner on SVG. In (a) multi-character example, DreamRunner generates videos with significantly better character consistency compared to other strong baselines, while others fail to maintain object consistency (e.g., VLogger), or fail to generate multiple objects ((a) Row 2,4). In (b) single-character setting, integrating SR3AI and locally-injected motion/subject priors consistently improve overall quality, complex motion synthesis and multi-event coherent composition.
@@ 4.2 Story-To-Video Generation Evaluation
Table 2: Ablation studies for the effectiveness of RAG and SR3AI in DreamRunner. Default DreamRunner achieves the best text-following ability and event transition smoothness.
@@ 4.3 Ablation Studies
In this section, we demonstrate the effectiveness of RAG for automatic video retrieval in motion prior learning and SR3AI to achieve fine-grained control over objects and their motion for coherent composition. As shown in Table 2, leveraging SR3AI for enhanced multi-object multi-event binding (2nd row) significantly improves the smooth transition between events within a single scene, while also enhancing visual quality and text following—potentially because decomposing attention into region-based components helps the model “divide and conquer” each part more effectively. Besides, incorporating retrieval-augmented motion prior learning (3rd row) boosts the video-text similarity both for fine-grained and full-prompt text following, showing its importance for motion enhancement. Lastly, we show that RAG and SR3AI can be combined effectively (last row), with the combined model achieving best results in smooth event transition, text alignment, and visual quality.
@@ 4.4 Qualitative Ablations and Comparisons
We provide qualitative ablations of our methods and comparison to previous SoTA in Fig. 4.
In Fig. 4 (a), we compare our method with SoTA and analyze the impact of our region-based attention and LoRA injection design for multi-character generation. Our approach significantly outperforms Vlogger [81] (rows 1 vs. 5), producing coherently composed characters and motions. Row 2 presents results from CogVideoX-2B using scene descriptions as conditions while injecting both character LoRAs globally. The outcome shows that global LoRA injection causes interference, resulting in a robot-like teddy bear rather than distinct characters, showing the importance of region-based LoRA injection for multi-character customization. To examine the effect of region-based attention design, we ablate it by using either hard-regional attention or full attention between visual latents (in Fig. 3 bottom-right part) in rows 4 and 5. Hard-regional attention (row 4) enforces strict adherence to the LLM layout plan (row 3) but prevents smooth transitions due to the lack of inter-region latent interaction. In contrast, full attention (row 5) enables smooth spatial and temporal transitions while ensuring that each region remains conditioned on its corresponding textual embeddings, resulting in high-quality multi-character, multi-motion composition and customization.
Model Consist-attr Dynamic-attr Spatial Motion Action Interaction
Gen-3 [5] 0.7045 0.2078 0.5533 0.3111 0.6280 0.7900
Dreamina [2] 0.8220 0.2114 0.6083 0.2391 0.6660 0.8175
PixVerse [3] 0.7370 0.1738 0.5874 0.2178 0.6960 0.8275
Kling [6] 0.8045 0.2256 0.6150 0.2448 0.6460 0.8475
VideoCrafter2 [10] 0.6750 0.1850 0.4891 0.2233 0.5800 0.7600
Open-Sora 1.2 [21] 0.6600 0.1714 0.5406 0.2388 0.5717 0.7400
Open-Sora-Plan v1.1.0 [31] 0.7413 0.1770 0.5587 0.2187 0.6780 0.7275
VideoTetris [55] 0.7125 0.2066 0.5148 0.2204 0.5280 0.7600
LVD [35] 0.5595 0.1499 0.5469 0.2699 0.4960 0.6100
CogVideoX-2B [68] 0.6775 0.2118 0.4848 0.2379 0.5700 0.7250
CogVideoX-2B+SR3A (Ours) 0.7350 (+8.5%) 0.2672 (+26.2%) 0.6040 (+24.6%) 0.2608 (+9.6%) 0.5840 (+2.5%) 0.7625 (+5.2%)
CogVideoX-5B [68] 0.7232 0.2250 0.5845 0.2551 0.6040 0.7995
CogVideoX-5B+SR3A (Ours) 0.7650 (+5.8%) 0.2832 (+25.9%) 0.6775 (+15.9%) 0.3041 (+19.2%) 0.6340 (+5.0%) 0.8725 (+9.1%)
We conduct detailed ablations of SR3AI and prior injections under the single-character setting in Fig. 4 (b). Directly using the base CogVideoX-2B model conditioned solely on the scene description without additional constraints (row 1) results in a vague character and background, along with abrupt scene transitions where the mermaid is absent in the second scene. This demonstrates that relying only on complex textual conditions in the base model leads to limited visual quality and fidelity. Row 2 presents results where only the character LoRA is injected globally into the model from row 1. This improves character consistency, allowing the mermaid to better follow the reference images. However, sudden scene transitions persist due to the complexity of the textual prompt, and the mermaid and the “inspecting” motion remains absent in the second scene. Moreover, global LoRA injection affects the background, making the kelp forest appear more cartoon-like—rounder and simplified—whereas in row 1, it retains a more natural, interwoven, and angular structure. Row 4 shows the results of integrating SR3AI with layout plan from row 3, supporting smooth transitions and well-followed character trajectories. The kelp forest appears more photorealistic thanks to locally injected character prior preventing interference with the background. However, the mermaid’s motion remains limited, particularly for the challenging “inspecting” action, due to the base model’s limitations in complex motion generation. This issue is largely solved when injecting RAG-learned motion priors in row 5. The mermaid exhibits significantly clearer motions: wanders from right to left, then stoops and moves her hand to scrabble on the kelp strand with shells, fully capturing the complex “inspecting” motion, thanks to the learned strong motion priors. Our full model generates high-fidelity videos adhere to the complex scene description, highlighting the importance of SR3AI for coherent composition of multiple objects/events, and our retrieval-augmented motion prior learning stage and injection strategy for complex motion synthesis.
@@
Table 4: Effect of RAG and per-caption prompt for motion prior learning. Both design improves motion synthesis quality.
@@ 4.5 Compositional T2V Generalization
In this section, we demonstrate how our spatial-temporal region-based attention module (SR3A) enhances compositional text-to-video generation, as evaluated on T2V-CompBench [54]. Here, we use only the attention module without LoRA injection, as no customization is required. Given a prompt, we employ an LLM (GPT-4o) to generate fine-grained layout plans, while SR3A ensures coherent composition of objects and events. As shown in Table 3, SR3A significantly improves CogVideoX-2B and CogVideoX-5B [68] across all categories. Notably, it boosts dynamic attribute binding (where an object’s attributes change over time) by 25+%, enhances spatial binding (ensuring multiple objects maintain predefined spatial relationships) by 15+%, and improves motion binding (objects moving in specified directions) by at least 10%. These results highlight SR3A’s ability to maintain coherent compositions involving multiple objects, trajectories, and sequential events. Additionally, SR3A significantly enhances fine-grained details such as attribute consistency, action binding, and object interactions, demonstrating its strong control capabilities. Beyond baseline comparisons, DreamRunner built on CogVideoX-5B achieves state-of-the-art performance in five dimensions compared to other open-source models and even outperforms all closed-source models in dynamic attribute binding, spatial binding, and object interactions. These results underscore our method’s effectiveness in advancing open-source models to match or exceed closed-source benchmarks, highlighting its adaptability to even more robust models. We provide 4 qualitative examples per dimension against baselines in Appendix D.
@@ 4.6 Effect of RAG and Per-Caption Prompt
We investigate the effectiveness of retrieval-augmented test-time fine-tuning and our per-caption prompt design for learning an enhanced motion prior. Specifically, for each motion, we use an LLM [41] to provide 6 prompts in the whole 64-motion set, then evaluate the average CLIP/ViCLIP scores using these prompts. As shown in Table 4, applying our approach to CogVideoX-2B results in both CLIP and ViCLIP score improvements. The substantial ViCLIP gain indicates stronger alignment between the story and the generated video, highlighting significant enhancements in motion accuracy. The row2-3 results also indicate that per-video prompts outperform single prompts, suggesting that video-specific prompts help the model implicitly disregard unrelated visual elements and focus on learning more motion-specific patterns. These results validate that RAG effectively retrieves videos with relevant motions, aiding in learning a more accurate motion prior for the model.
@@ A.1 Retrieval Processes
    Initial Retrieval with BM25: We use the text-only BM25 score [49] based on video captions in the dataset to retrieve 400 candidate videos for the query. To ensure the retrieved videos are human motion-centric, we add "person is" at the beginning of the query ("person is sitting").
@@
    Clip Segmentation via Object Tracking: We track individuals within videos using YOLOv5 [26] and segment clips into human-centered segments based on the tracking results, keeping meaningful human-focused content.
@@
    Scoring with CLIP and ViCLIP [47, 57]: To ensure the fidelity between the segmented video clips and the query, we compute semantic similarity scores to the query text (e.g., “the person is sitting”) using CLIP and ViCLIP for each segmented clip. The CLIP score is computed by sampling eight frames and averaging frame-query scores, while the ViCLIP score is directly computed on the full video and query. We select the top 20 videos that satisfy the average scores of CLIP and ViCLIP > 0.2. We retain the top four videos based on their ranking if fewer than four videos meet this threshold.

### v4

@@ 3.3 Spatial-Temporal-Region-Based Diffusion
We build our model on CogVideoX (yang2024cogvideox), a text-to-video generation model designed on top of a Diffusion Transformer (DiT). Unlike methods that use separate spatial and temporal attention for efficient video modeling, CogVideoX employs a 3D full attention module, integrating self-attention across concatenated embeddings of all visual latents and the text condition embeddings. We extend this module to enable region-specific conditioning via masking, aligning different regions with their respective text descriptions. Specifically, given a fine-grained plan with $N$ region-specific text descriptions $C_{1},C_{2},\dots,C_{N}$ and corresponding layouts $L_{1},L_{2},\dots,L_{N}$ across frames, we encode each text condition $C_{i}$ to produce embeddings $T_{1},T_{2},\dots,T_{N}$ (Figure 2 top right). At each attention layer, we identify the visual latents corresponding to each layout $L_{i}$ in the latent space. We then perform masked self-attention on the concatenation of $T_{1},T_{2},\dots,T_{N}$ and $L_{1},L_{2},\dots,L_{N}$. The self-attention mask is defined as follows: for each region’s visual latents $L_{i}$, attention is allowed to its corresponding text condition embeddings $T_{i}$ and all visual latents $L_{1},L_{2},\dots,L_{N}$. Conversely, for each condition embeddings $T_{i}$, attention is restricted to itself and its corresponding latents $L_{i}$. This design ensures each region is conditioned on its specific textual description while maintaining interactions among visual latents through unmasked attention among $L_{1},L_{2},\dots,L_{N}$. No modifications are made to other modules in the base model, preserving the integrity of its original architecture. A visualization example of such masking strategy is contained in Fig. 3.
@@
Model Consist-attr Dynamic-attr Spatial Motion Action Interaction
Gen-3 (gen3) 0.7045 0.2078 0.5533 0.3111 0.6280 0.7900
Dreamina (Dreamina) 0.8220 0.2114 0.6083 0.2391 0.6660 0.8175
PixVerse (PixVerse) 0.7370 0.1738 0.5874 0.2178 0.6960 0.8275
Kling (kling) 0.8045 0.2256 0.6150 0.2448 0.6460 0.8475
VideoCrafter2 (chen2024videocrafter2) 0.6750 0.1850 0.4891 0.2233 0.5800 0.7600
Open-Sora 1.2 (opensora) 0.6600 0.1714 0.5406 0.2388 0.5717 0.7400
Open-Sora-Plan v1.1.0 (pku_yuan_lab_and_tuzhan_ai_etc_2024_10948109) 0.7413 0.1770 0.5587 0.2187 0.6780 0.7275
VideoTetris (videoteris) 0.7125 0.2066 0.5148 0.2204 0.5280 0.7600
LVD (lian2023llm) 0.5595 0.1499 0.5469 0.2699 0.4960 0.6100
CogVideoX-2B (yang2024cogvideox) 0.6775 0.2118 0.4848 0.2379 0.5700 0.7250
CogVideoX-2B+SR3A (Ours) 0.7350 (+8.5%) 0.2672 (+26.2%) 0.6123 (+26.3%) 0.2608 (+9.6%) 0.5840 (+2.5%) 0.7625 (+5.2%)
CogVideoX-5B (yang2024cogvideox) 0.7232 0.2250 0.5845 0.2551 0.6040 0.7995
CogVideoX-5B+SR3A (Ours) 0.7650 (+5.8%) 0.2832 (+25.9%) 0.6875 (+17.5%) 0.3041 (+19.2%) 0.6340 (+5.0%) 0.8725 (+9.1%)
@@
Table 3: Ablation studies for the effectiveness of RAG and SR3AI in DreamRunner. Our full model achieves the best text-following ability and event transition smoothness.
4 Experiments
In this section, we first introduce the evaluation datasets and evaluation metrics details in Section 4.1, then compare our DreamRunner with prior methods on story-to-video generation in Section 4.2. Next, we present detailed ablation studies on the necessity of RAG and effectiveness of SR3AI in Section 4.3, and demonstrate the generalizability of our DreamRunner to improve compositional text-to-video generation on T2V-CompBench (sun2024t2v) in Section 4.5. Moreover, we show the effectiveness of RAG for learning the motion prior on a more comprehensive motion dataset we collect in Section 4.6. Lastly, we present qualitative comparison between our DreamRunner and previous approaches in Section 4.4.
4.1 Experimental Setups
Evaluation Datasets. We evaluate DreamRunner on two tasks: (1) story-to-video generation, and (2) compositional text-to-video generation. The first task focuses on the model’s ability to follow the text closely while maintaining character and scene consistency throughout the story. The second task assesses various aspects of compositionality in video generation. For (1) story-to-video generation, we collect and introduce a new benchmark dataset, DreamStorySet. Specifically, we collect 10 characters, including 6 from existing customization datasets (CustomConcept101 (kumari2023multi) and Dreambooth (ruiz2023dreambooth)), and 4 with generation models (FLUX (flux)). (featuring two motions per scene) and three multi-character stories (featuring two or three motions per scene). Each story comprises 5 to 8 scenes, incorporating a total of 64 diverse motions throughout. We focus on single-character stories for quantitative evaluation of SVG models and reserve multi-character stories for qualitative evaluation. For (2) compositional text-to-video generation, we use the T2V-CompBench (sun2024t2v) to benchmark the performance of DreamRunner, where we select six dimensions except numeracy.
Evaluation Metrics. We evaluate our storytelling videos across multiple dimensions: Character Consistency (Frame-to-Reference CLIP/DINO), Text Alignment at both the full narration and fine-grained scene level (Image/Video-to-Text CLIP/ViCLIP), and Transition Smoothness (Frame-to-Frame DINO). For visual quality, we adopt three representative metrics from VBench (huang2023vbench): aesthetic quality, imaging quality, and video smoothness (li2023amt), with full quality results and metric details provided in the Appendix. We follow similar evaluation metrics to T2V-ComBench (sun2024t2v) for compositional T2V.
4.2 Story-To-Video Generation Evaluation
We compare DreamRunner with prior SoTAs (VideoDirectorGPT (lin2023videodirectorgpt) and VLogger (zhuang2024vlogger)) on our DreamStorySet dataset for story-to-video generation. For fairness, each scene narration is split into two single-motion descriptions, with corresponding videos later merged into a single-scene video. As shown in Table 1, DreamRunner improves CLIP/DINO scores by 13.1%/33.4% over VLogger, demonstrating the effectiveness of our learned subject prior and region-based LoRA injection for character consistency. To evaluate text-following capability, we assess both full-prompt adherence and fine-grained event alignment. DreamRunner improves CLIP/ViCLIP scores consistently on both settings, showing superior alignment with both full-scene and fine-grained event descriptions. For transition quality, we compute the DINO-based transition score to measure scene and event consistency. DreamRunner improves transitions by 27.2% over VLogger, highlighting the effectiveness of SR3AI in generating sequential events in a single scene. Lastly, we evaluate visual quality across aesthetic quality, imaging quality, and motion smoothness. DreamRunner enhances aesthetics while slightly improving the other two, demonstrating its capability to generate high-quality videos adhere to complex scene descriptions with smooth event transitions. We provide three additional quality scores from VBench and qualitative examples in the Appendix.
Figure 4: Qualitative comparison and ablations of DreamRunner on SVG. In (a) multi-character example, DreamRunner produces significantly better character consistency compared to other strong baselines, while others fail to maintain object consistency (e.g., VLogger), or fail to generate multiple objects ((a) Row 2,4). In (b) single-character setting, integrating SR3AI and locally-injected priors consistently improve overall quality, complex motion synthesis and coherent composition. Note that in the overlapped regions in (b) row 3, the caption is a merge of the two. For cleaner visualization, we don’t show it here.
@@ 4.3 Ablation Studies
In this section, we demonstrate the effectiveness of RAG for automatic video retrieval in motion prior learning and SR3AI for fine-grained control over objects and their motions to achieve coherent composition. As shown in Table 3, using SR3AI for enhanced multi-object multi-event binding (2nd row) significantly improves event transition smoothness within a single scene, as well as visual quality and text alignment—likely because decomposing attention into region-based components enables the model to “divide and conquer” more effectively. Incorporating retrieval-augmented motion prior learning (3rd row) further improves video-text similarity for both fine-grained and full-prompt alignment, showing its importance for motion enhancement. Finally, combining both(last row) yields the best results across event transition, text alignment, and visual quality. We also provide additional ablations on RAG pipeline, Layer-seperation strategy for motion prior learning, and computational cost comparison in Appendix B.
@@ 4.4 Qualitative Ablations and Comparisons
We provide qualitative comparisons and ablations in Fig. 4.
In Fig. 4(a), we compare DreamRunner with VLogger(zhuang2024vlogger) for multi-character generation and analyze the effects of region-based attention and LoRA injection. Our method (row 5) generates coherent multi-character composition and motions, outperforming VLogger (row 1). Using CogVideoX-2B with character LoRAs injected globally (row 2) results in interference, producing a robot-like teddy bear and blurry compositions, highlighting the necessity of region-based LoRA injection. To study attention design, we ablate our region-based attention by comparing hard-regional attention (row 4) and full attention (row 5). While hard-regional attention strictly follows the layout plan (row 3), it limits spatial-temporal continuity due to lack of inter-region interaction. In contrast, our full attention mechanism enables smooth transitions while maintaining spatial-region constraints, supporting high-quality multi-character customization.
In Fig. 4(b), we present single-character ablations. Using the base CogVideoX-2B with only scene-level text (row 1) leads to vague character/background and missing actions. Injecting global character LoRA (row 2) improves character appearance but still fails on action and transition quality, and degrades background fidelity (e.g., cartoonish kelp forest). Applying SR3AI with layout plans (row 4) improves trajectory control and preserves background fidelity through localized injection, but motion remains limited. Injecting RAG-learned motion priors (row 5) enables clear, fine-grained motion execution (e.g., the mermaid stooping and interacting with kelp), demonstrating the benefit of our motion prior learning and injection strategy. Overall, our full model combines coherent composition with strong motion quality, showing the effectiveness of SR3AI and retrieval-augmented prior learning for complex, multi-entity video generation.
@@
Table 4: Effect of RAG and per-video prompt for motion prior learning.
@@ 4.5 Compositional T2V Generalization
In this section, we demonstrate how our spatial-temporal region-based attention module (SR3A) enhances compositional T2V, as evaluated on T2V-CompBench (sun2024t2v). We use SR3A (no LoRA injection) as no customization is required. Given a prompt, we use GPT-4o to generate layout plans, and SR3A ensures coherent composition of objects and events. As shown in Table 2, SR3A significantly improves both CogVideoX-2B and CogVideoX-5B (yang2024cogvideox) across all categories. Specifically, it boosts dynamic attribute binding by over 25%, spatial binding by over 15%, and motion binding by at least 10%, highlighting SR3A’s ability to maintain coherent multi-object compositions, trajectories, and sequential events. It also improves scores on other fine-grained aspects, demonstrating strong control capabilities. Notably, DreamRunner built on CogVideoX-5B achieves SoTA results in five dimensions among open-source models, and surpasses all closed-source models in dynamic attribute binding, spatial binding, and object interactions, highlighting its ability to close the open-closed source model gap and adapt to stronger base models. Qualitative examples per dimension are in Appendix E.
@@ 4.6 Effect of RAG and Per-Caption Prompt
We investigate the effectiveness of retrieval-augmented test-time fine-tuning and our per-caption prompt design for learning an enhanced motion prior. Specifically, for each motion in the 64-motion set, we use GPT-4o to generate six prompts and evaluate the average CLIP/ViCLIP scores. As shown in Table 4, applying our approach to CogVideoX-2B improves both scores, with the significant ViCLIP gain indicating better story-video alignment and enhanced motion accuracy. Rows 2–3 further show that per-video prompts outperform single prompts, suggesting that video-specific conditioning helps the model ignore unrelated visual cues and better capture motion-specific patterns. These results confirm that RAG effectively retrieves motion-relevant videos and facilitates the learning of more accurate motion priors.
@@ A.1 Retrieval Processes
    Initial Retrieval with BM25: We use the text-only BM25 score (robertson2009probabilistic) based on video captions in the dataset to retrieve 400 candidate videos for the query. To ensure the retrieved videos are human motion-centric, we add "person is" at the beginning of the query ("person is sitting").
@@
    Clip Segmentation via Object Tracking: We track individuals within videos using YOLOv5 (Jocher_YOLOv5_by_Ultralytics_2020) and segment clips into human-centered segments based on the tracking results, keeping meaningful human-focused content.
@@
    Scoring with CLIP and ViCLIP (radford2021learning; wang2022internvideo): To ensure the fidelity between the segmented video clips and the query, we compute semantic similarity scores to the query text (e.g. “the person is sitting”) using CLIP and ViCLIP for each segmented clip. The CLIP score is computed by sampling eight frames and averaging frame-query scores, while the ViCLIP score is directly computed on the full video and query. We select the top 20 videos that satisfy the average scores of CLIP and ViCLIP > 0.2. We retain the top four videos based on their ranking if fewer than four videos meet this threshold.